# Generalized Spherical Neural Operators: Green’s Function Formulation

### v1

@@ A.1 Properties of spherical harmonics
The explicit form of spherical harmonic function (muller2006spherical) is:
@@
Spherical harmonic functions satisfy the orthonormality condition (muller2006spherical):
@@
For subsequent analysis, we recall how the spherical harmonics transform under a rotation operation $R$ from $SO(3)$. The rotation of the spherical harmonic function can be expressed as (driscoll1994computing):
@@
where $D_{k,m}^{(l)}$ denotes the Wigner D-matrix (edmonds1996angular) and $\Lambda(R)$ represents the rotation operator acting on spherical functions. Taking the complex conjugate of Equation equation 23 yields:
@@ A.3 Orthogonality condition screening non-zero terms
Parameterize $R\in SO(3)$ using Euler angles $R=u(\phi)a(\theta)u(\psi)$ (driscoll1994computing), with inverse $R^{-1}=u(-\psi)a(-\theta)u(-\phi)$. The corresponding D-matrix is:
@@
which d is the Wigner d-matrix (edmonds1996angular). Then Taking the complex conjugate:
@@
For any rotation $u(\psi)$ on the z-axis, substitute $R\to Ru(\psi)$. Due to right-invariance of the Haar measure $dR$ (driscoll1994computing), the remaining integral is adjusted to:
@@ A.4 Relationship between D-matrices and spherical harmonics
Because the unitary characteristic of the Wigner D-matrix (driscoll1994computing), we can get the relationship between D-matrix and spherical Harmonic function:
@@ B.2 Generalized Spherical Harmonic Network
For WB and SWE experiments, this study uses Generalized Spherical Fourier Network (GSHNet) with a depth of 2, which performs two spherical downsampling operations. Specifically, the encoder first expands the input channel to $C$. In the first two GSNO blocks, the spatial resolution, i.e., the number of latitudinal and longitudinal sampling points in the inverse spherical harmonic transform (ISHT), is reduced by half, while the channel dimension is doubled via MLPs and convolutional layers.
Subsequently, the next two GSNO blocks perform upsampling by doubling the number of samples in latitude and longitude while continuing to transform the channels. The final GSNO block is used as the output layer: it projects the features back to channel dimension $C$ without applying further spatial scaling. The implementations of MLPs, encoders and decoders are consistent with those used in SFNONet (bonev2023spherical). The kernel of all convolution operations is $1\times 1$ for channel adjustment. GSHNet also includes three skip connections to enhance gradient flow and multi-scale feature fusion: The output of the final GSNO block is concatenated with the original input; The output of the fourth GSNO block is added to the input of the first GSNO block; The output of the third GSNO block is added to the input of the second GSNO block.
@@ C.1 Model implementation detail
The key model parameter settings used for the two experiments are presented in Table 5. Our proposed GSHNet with a depth of 2 consists of five blocks, including two downsampling layers, two upsampling layers, and an output layer. Furthermore, the embedding dimension of GSHNet is set to a smaller value (8 and 64), as the depth increases, the embedding dimension gradually increases until it is four times the original dimension. The depth of other models represents the number of core blocks.
@@
Hyperparameters
ClimaX FourCastNet SFNONet GSHNet
@@
The loss function $\mathcal{L}$ of the proposed method is the spherical grid weighted mean relative error between the prediction $F(X_{n})$ and the target $Y_{n+t}$, calculated following Bonus et al. (bonev2023spherical) as follows:
@@
where $v_{i,j}$ is the products of the Jacobian $sin(\lambda_{i})$ ($\lambda_{i}$ represents the latitude at grid point) and the quadrature weights (bonev2023spherical). $C$ is the number of predicted variables. $n$ is the index of the initial time step, $t$ is the predicted autoregressive steps. This loss function is used in all the experiments.
@@ C.2 Spherical shallow water equations
Spherical Shallow Water Equations (SSWE) form a nonlinear hyperbolic PDE system that models the motion of thin-layer fluids on a rotating sphere. The core underlying assumption is the shallow water approximation, where the vertical scale of the fluid layer is much smaller than the horizontal scale (bonev2018discontinuous).
The dataset is publicly available, which can be obtained and used for evaluation through Torch-Harmonics GitHub  (bonev2023spherical). Consistent with Bonus et al. (bonev2023spherical), we use mean relative error of each variable as the evaluation metric.
@@
Table 6: Supplementary experiments on SWE at $128\times 256$ resolution (MRE↓) and WB at $64\times 128$ resolution (ACC↑).
Methods
SWE at 5h
SWE at 10h
WB at 1day
WB at 3day
WB at 5day
@@ C.3 Weather forecasting
The dataset is publicly available at WeatherBench GitHub (rasp2020weatherbench). Following (liu2024evaluation), the chosen 24 common variables are detailed in Table 9, including 10U, 10V, 2T, U50, U50, U250, U500, U600, U700, U850, U925, V50, V250, V500, V600, V700, V850, V925, T50, T250, T500, T600, T700, T850, T925. And 10U, 10V, 2T, U600, V600 and T600 are prediction target.
@@ 1 Day Forecast
Climax (nguyen2023climax) - 4.23 ± 0.43 80.9 ± 5.16 113 ± 7.04 67.9 ± 4.17 118 ± 7.71 6.68 ± 0.51 66.8 ± 2.06
FourCastNet (pathak2022fourcastnet) - 4.36 ± 0.45 94.2 ± 5.95 132 ± 8.18 78.9 ± 5.17 137 ± 9.30 7.98 ± 0.66 75.7 ± 2.45
SFNONet (bonev2023spherical) SFNO 5.18 ± 0.58 175 ± 12.1 242 ± 17.0 134 ± 8.98 234 ± 16.7 10.8 ± 0.93 134 ± 4.70
GSHNet (Ours) GSFNO 4.22 ± 0.40 124 ± 8.72 173 ± 10.4 106 ± 7.18 168 ± 11.3 8.75 ± 0.81 97.5 ± 3.18
SFNONet (bonev2023spherical) GSFNO 4.82 ± 0.52 162 ± 10.5 221 ± 14.9 127 ± 8.32 215 ± 14.8 9.97 ± 0.87 123 ± 4.16
GSHNet SFNO 4.76 ± 0.48 156 ± 9.90 208 ± 14.1 122 ± 8.34 204 ± 14.1 9.89 ± 0.86 117 ± 3.97
@@ 3 Day Forecast
Climax (nguyen2023climax) - 11.6 ± 1.55 384 ± 35.0 545 ± 56.1 305 ± 28.8 558 ± 61.4 31.4 ± 4.38 306 ± 15.8
FourCastNet (pathak2022fourcastnet) - 12.7 ± 1.66 402 ± 33.1 594 ± 54.4 322 ± 28.0 616 ± 62.3 33.9 ± 4.51 330 ± 15.6
SFNONet (bonev2023spherical) SFNO 10.3 ± 1.51 372 ± 32.0 535 ± 51.0 295 ± 26.7 546 ± 60.5 30.3 ± 4.34 298 ± 14.9
GSHNet (Ours) GSFNO 9.08 ± 1.32 335 ± 26.9 478 ± 44.7 267 ± 24.0 492 ± 51.4 26.7 ± 3.79 268 ± 13.9
SFNONet (bonev2023spherical) GSFNO 9.59 ± 1.45 357 ± 29.3 501 ± 48.1 281 ± 25.5 524 ± 56.2 28.5 ± 4.10 284 ± 14.0
GSHNet SFNO 9.50 ± 1.40 362 ± 29.3 495 ± 47.7 284 ± 25.9 522 ± 55.8 28.1 ± 3.91 282 ± 13.9
@@ 5 Day Forecast
Climax (nguyen2023climax) - 18.0 ± 3.15 572 ± 53.6 831 ± 91.2 518 ± 49.4 922 ± 121 57.7 ± 9.46 486 ± 28.1
FourCastNet (pathak2022fourcastnet) - 20.9 ± 3.18 510 ± 41.7 725 ± 70.7 471 ± 39.3 811 ± 99.5 58.9 ± 8.24 433 ± 22.5
SFNONet (bonev2023spherical) SFNO 14.4 ± 2.57 502 ± 46.5 726 ± 81.6 440 ± 42.9 794 ± 105 48.1 ± 7.76 421 ± 24.6
GSHNet (Ours) GSFNO 13.3 ± 2.34 471 ± 42.2 687 ± 73.4 411 ± 40.5 762 ± 93.8 45.9 ± 7.32 398 ± 22.2
SFNONet (bonev2023spherical) GSFNO 13.9 ± 2.46 484 ± 44.0 699 ± 75.1 427 ± 41.4 776 ± 94.9 47.1 ± 7.57 408 ± 22.6
GSHNet SFNO 13.8 ± 2.45 492 ± 44.8 705 ± 76.3 430 ± 41.5 783 ± 96.0 46.8 ± 7.54 412 ± 22.9
@@
The anomaly correlation coefficient (ACC) and the latitude-weighted mean square error (MSE) are used to measure the performance of the model in WeatherBench (WB). MSE between the prediction $F(X_{n})$ and the target $Y_{n+t}$ for evaluation is calculated following Rasp et al. (nguyen2023climax) as:
@@
where $H$ is the number of latitudes and $W$ is the number of longitudes. The latitude weight (nguyen2023climax) are calculated as:
@@
ACC (nguyen2023climax) is used to measure spatial correlation between predicted anomalies $F(X_{n})^{\prime}$ and target anomalies $Y_{n+t}^{\prime}$:
@@
The results in the Table 3 clearly demonstrate that our spectral implementation (GSNO) consistently and significantly outperforms the SFNO and spatial position-encoding SNO across all forecast lead times. This indicates that the spectral implementation derived from our Green’s function framework is, in itself, a more effective and different design. Moreover, the spatial position encoding added in each operator ($SHT[f+C_{f}\cdot[G]_{\theta(x,y)}]$) is not conducive to long-term stable prediction.
@@ C.4 Diffusion MRI modeling of brain microstructure
We randomly select dMRI images of 30 subjects in the Human Connectome Project (HCP) (van2013wu), where 20 subjects were for training, 5 for validation, and 5 for test. Detailed preprocessing. The complete dMRI data of each subject contains 288 volumes (multi-shell HARDI), including 270 volumes of b=1000, 2000, 3000 $s/mm^{2}$ with 90 gradient directions for each shell and 18 b0 volumes. Due to the wide application of low b-value with 32 gradient directions in clinical practice (zeng2022fod), we subsample 32 volumes of b=1000 and 1 volume of b0 from the complete dMRI according to the HCP protocol, to obtain single-shell LARDI. Further, we use MSMT-CSD (jeurissen2014multi) on the multi-shell HARDI to obtain high angular resolution fiber orientation distribution (FOD) as the ground truth (HAR-FOD). Meanwhile, we use SSMT-CSD (khan2020three) on the single-shell LARDI to obtain single-shell low angular resolution FOD as the condition of GSNO (LAR-FOD). $l_{max}$ is set to the default value of 8 to balance precision and complexity (zeng2022fod).
SSMT-CSD (khan2020three), FOD-Net (zeng2022fod),  FOD-SFNO (bonev2023spherical), ESCNN (snoussi2025equivariant) and our proposed FOD-GSNOadopt the same network architecture as FOD-Net. The difference lies in that FOD-Net uses 3D convolutional layers to handle a large number of voxels, while FOD-SFNO, ESCNN and FOD-GSNO stack voxels on the channel dimension and then use the corresponding operators for processing. To ensure a fair comparison, FOD-SFNO and FOD-GSNO have exactly the same hyperparameters, with the only difference being the types of operators. And their parameter comparison is shown in Table 10.
@@ C.6 Detailed computational trade-offs
We add the comparison of parameters, FLOPS, training time (seconds per epoch) and inference time for two models, SFNONet and GSHNet. All results are obtained using the same hardware and dataset (Weatherbench).
@@
Models GSHNet (SFNO) GSHNet (GSNO) SFNONet (SFNO) SFNONet (GSNO)

### v2

@@ A.1 Properties of spherical harmonics
The explicit form of spherical harmonic function (Müller, 2006) is:
@@
Spherical harmonic functions satisfy the orthonormality condition (Müller, 2006):
@@
For subsequent analysis, we recall how the spherical harmonics transform under a rotation operation $R$ from $SO(3)$. The rotation of the spherical harmonic function can be expressed as (Driscoll and Healy, 1994):
@@
where $D_{k,m}^{(l)}$ denotes the Wigner D-matrix (Edmonds, 1996) and $\Lambda(R)$ represents the rotation operator acting on spherical functions. Taking the complex conjugate of Equation equation 23 yields:
@@ A.3 Orthogonality condition screening non-zero terms
Parameterize $R\in SO(3)$ using Euler angles $R=u(\phi)a(\theta)u(\psi)$ (Driscoll and Healy, 1994), with inverse $R^{-1}=u(-\psi)a(-\theta)u(-\phi)$. The corresponding D-matrix is:
@@
which d is the Wigner d-matrix (Edmonds, 1996). Then Taking the complex conjugate:
@@
For any rotation $u(\psi)$ on the z-axis, substitute $R\to Ru(\psi)$. Due to right-invariance of the Haar measure $dR$ (Driscoll and Healy, 1994), the remaining integral is adjusted to:
@@ A.4 Relationship between D-matrices and spherical harmonics
Because the unitary characteristic of the Wigner D-matrix (Driscoll and Healy, 1994), we can get the relationship between D-matrix and spherical Harmonic function:
@@ B.2 Generalized Spherical Harmonic Network
For WB and SSWE experiments, this study uses Generalized Spherical Fourier Network (SHNet) with a depth of 2, which performs two spherical downsampling operations. Specifically, the encoder first expands the input channel to $C$. In the first two GSNO blocks, the spatial resolution, i.e., the number of latitudinal and longitudinal sampling points in the inverse spherical harmonic transform (ISHT), is reduced by half, while the channel dimension is doubled via MLPs and convolutional layers.
Subsequently, the next two GSNO blocks perform upsampling by doubling the number of samples in latitude and longitude while continuing to transform the channels. The final GSNO block is used as the output layer: it projects the features back to channel dimension $C$ without applying further spatial scaling. The implementations of MLPs, encoders and decoders are consistent with those used in SFNONet (Bonev et al., 2023). The kernel of all convolution operations is $1\times 1$ for channel adjustment. SHNet also includes three skip connections to enhance gradient flow and multi-scale feature fusion: The output of the final GSNO block is concatenated with the original input; The output of the fourth GSNO block is added to the input of the first GSNO block; The output of the third GSNO block is added to the input of the second GSNO block.
@@ C.1 Model implementation detail
The key model parameter settings used for the two experiments are presented in Table 5. Our proposed SHNet with a depth of 2 consists of five blocks, including two downsampling layers, two upsampling layers, and an output layer. Furthermore, the embedding dimension of SHNet is set to a smaller value (8 and 64), as the depth increases, the embedding dimension gradually increases until it is four times the original dimension. The depth of other models represents the number of core blocks.
@@
Hyperparameters ClimaX FourCastNet SFNONet SHNet
@@
The loss function $\mathcal{L}$ of the proposed method is the spherical grid weighted mean relative error between the prediction $F(X_{n})$ and the target $Y_{n+t}$, calculated following Bonus et al. (Bonev et al., 2023) as follows:
@@
where $v_{i,j}$ is the products of the Jacobian $sin(\lambda_{i})$ ($\lambda_{i}$ represents the latitude at grid point) and the quadrature weights (Bonev et al., 2023). $C$ is the number of predicted variables. $n$ is the index of the initial time step, $t$ is the predicted autoregressive steps. This loss function is used in all the experiments.
@@ C.2 Spherical shallow water equations
Spherical Shallow Water Equations (SSWE) form a nonlinear hyperbolic PDE system that models the motion of thin-layer fluids on a rotating sphere. The core underlying assumption is the shallow water approximation, where the vertical scale of the fluid layer is much smaller than the horizontal scale (Bonev et al., 2018).
The dataset is publicly available, which can be obtained and used for evaluation through Torch-Harmonics GitHub  (Bonev et al., 2023). Consistent with Bonus et al. (Bonev et al., 2023), we use mean relative error of each variable as the evaluation metric.
@@
Table 6: Supplementary experiments on SSWE at $128\times 256$ resolution (MRE↓) and WB at $64\times 128$ resolution (ACC↑).
Methods SSWE at 5h SSWE at 10h WB at 1day WB at 3day WB at 5day
@@ C.3 Weather forecasting
The dataset is publicly available at WeatherBench GitHub (Rasp et al., 2020). Following (Liu et al., 2024), the chosen 24 common variables are detailed in Table 9, including 10U, 10V, 2T, U50, U50, U250, U500, U600, U700, U850, U925, V50, V250, V500, V600, V700, V850, V925, T50, T250, T500, T600, T700, T850, T925. And 10U, 10V, 2T, U600, V600 and T600 are prediction target.
@@ 1 Day Forecast
Climax (Nguyen et al., 2023) - 4.23 ± 0.43 80.9 ± 5.16 113 ± 7.04 67.9 ± 4.17 118 ± 7.71 6.68 ± 0.51 66.8 ± 2.06
FourCastNet (Pathak et al., 2022) - 4.36 ± 0.45 94.2 ± 5.95 132 ± 8.18 78.9 ± 5.17 137 ± 9.30 7.98 ± 0.66 75.7 ± 2.45
SFNONet (Bonev et al., 2023) SFNO 5.18 ± 0.58 175 ± 12.1 242 ± 17.0 134 ± 8.98 234 ± 16.7 10.8 ± 0.93 134 ± 4.70
SHNet (Ours) GSNO 4.22 ± 0.40 124 ± 8.72 173 ± 10.4 106 ± 7.18 168 ± 11.3 8.75 ± 0.81 97.5 ± 3.18
SFNONet (Bonev et al., 2023) GSNO 4.82 ± 0.52 162 ± 10.5 221 ± 14.9 127 ± 8.32 215 ± 14.8 9.97 ± 0.87 123 ± 4.16
SHNet SFNO 4.76 ± 0.48 156 ± 9.90 208 ± 14.1 122 ± 8.34 204 ± 14.1 9.89 ± 0.86 117 ± 3.97
@@ 3 Day Forecast
Climax (Nguyen et al., 2023) - 11.6 ± 1.55 384 ± 35.0 545 ± 56.1 305 ± 28.8 558 ± 61.4 31.4 ± 4.38 306 ± 15.8
FourCastNet (Pathak et al., 2022) - 12.7 ± 1.66 402 ± 33.1 594 ± 54.4 322 ± 28.0 616 ± 62.3 33.9 ± 4.51 330 ± 15.6
SFNONet (Bonev et al., 2023) SFNO 10.3 ± 1.51 372 ± 32.0 535 ± 51.0 295 ± 26.7 546 ± 60.5 30.3 ± 4.34 298 ± 14.9
SHNet (Ours) GSNO 9.08 ± 1.32 335 ± 26.9 478 ± 44.7 267 ± 24.0 492 ± 51.4 26.7 ± 3.79 268 ± 13.9
SFNONet (Bonev et al., 2023) GSNO 9.59 ± 1.45 357 ± 29.3 501 ± 48.1 281 ± 25.5 524 ± 56.2 28.5 ± 4.10 284 ± 14.0
SHNet SFNO 9.50 ± 1.40 362 ± 29.3 495 ± 47.7 284 ± 25.9 522 ± 55.8 28.1 ± 3.91 282 ± 13.9
@@ 5 Day Forecast
Climax (Nguyen et al., 2023) - 18.0 ± 3.15 572 ± 53.6 831 ± 91.2 518 ± 49.4 922 ± 121 57.7 ± 9.46 486 ± 28.1
FourCastNet (Pathak et al., 2022) - 20.9 ± 3.18 510 ± 41.7 725 ± 70.7 471 ± 39.3 811 ± 99.5 58.9 ± 8.24 433 ± 22.5
SFNONet (Bonev et al., 2023) SFNO 14.4 ± 2.57 502 ± 46.5 726 ± 81.6 440 ± 42.9 794 ± 105 48.1 ± 7.76 421 ± 24.6
SHNet (Ours) GSNO 13.3 ± 2.34 471 ± 42.2 687 ± 73.4 411 ± 40.5 762 ± 93.8 45.9 ± 7.32 398 ± 22.2
SFNONet (Bonev et al., 2023) GSNO 13.9 ± 2.46 484 ± 44.0 699 ± 75.1 427 ± 41.4 776 ± 94.9 47.1 ± 7.57 408 ± 22.6
SHNet SFNO 13.8 ± 2.45 492 ± 44.8 705 ± 76.3 430 ± 41.5 783 ± 96.0 46.8 ± 7.54 412 ± 22.9
@@
The anomaly correlation coefficient (ACC) and the latitude-weighted mean square error (MSE) are used to measure the performance of the model in WeatherBench (WB). MSE between the prediction $F(X_{n})$ and the target $Y_{n+t}$ for evaluation is calculated following Rasp et al. (Nguyen et al., 2023) as:
@@
where $H$ is the number of latitudes and $W$ is the number of longitudes. The latitude weight (Nguyen et al., 2023) are calculated as:
@@
ACC (Nguyen et al., 2023) is used to measure spatial correlation between predicted anomalies $F(X_{n})^{\prime}$ and target anomalies $Y_{n+t}^{\prime}$:
@@
The results in the Table 3 clearly demonstrate that our spectral implementation (GSNO) consistently and significantly outperforms the SFNO and spatial position-embedding SNO across all forecast lead times. This indicates that the spectral implementation derived from our Green’s function framework is, in itself, a more effective and different design. Moreover, the spatial position embedding added in each operator ($SHT[f+C_{f}\cdot[G]_{\theta(x,y)}]$) is not conducive to long-term stable prediction.
@@ C.4 Diffusion MRI modeling of brain microstructure
We randomly select dMRI images of 30 subjects in the Human Connectome Project (HCP) (Van Essen et al., 2013), where 20 subjects were for training, 5 for validation, and 5 for test. Detailed preprocessing. The complete dMRI data of each subject contains 288 volumes (multi-shell HARDI), including 270 volumes of b=1000, 2000, 3000 $s/mm^{2}$ with 90 gradient directions for each shell and 18 b0 volumes. Due to the wide application of low b-value with 32 gradient directions in clinical practice (Zeng et al., 2022), we subsample 32 volumes of b=1000 and 1 volume of b0 from the complete dMRI according to the HCP protocol, to obtain single-shell LARDI. Further, we use MSMT-CSD (Jeurissen et al., 2014) on the multi-shell HARDI to obtain high angular resolution fiber orientation distribution (FOD) as the ground truth (HAR-FOD). Meanwhile, we use SSMT-CSD (Khan et al., 2020) on the single-shell LARDI to obtain single-shell low angular resolution FOD as the condition of GSNO (LAR-FOD). $l_{max}$ is set to the default value of 8 to balance precision and complexity (Zeng et al., 2022).
SSMT-CSD (Khan et al., 2020), FOD-Net (Zeng et al., 2022),  FOD-SFNO (Bonev et al., 2023), ESCNN (Snoussi and Karimi, 2025) and our proposed FOD-GSNOadopt the same network architecture as FOD-Net. The difference lies in that FOD-Net uses 3D convolutional layers to handle a large number of voxels, while FOD-SFNO, ESCNN and FOD-GSNO stack voxels on the channel dimension and then use the corresponding operators for processing. To ensure a fair comparison, FOD-SFNO and FOD-GSNO have exactly the same hyperparameters, with the only difference being the types of operators. And their parameter comparison is shown in Table 10.
@@ C.6 Detailed computational trade-offs
We add the comparison of parameters, FLOPS, training time (seconds per epoch) and inference time for two models, SFNONet and SHNet. All results are obtained using the same hardware and dataset (Weatherbench).
@@
Models SHNet (SFNO) SHNet (GSNO) SFNONet (SFNO) SFNONet (GSNO)